# Population Transformer: Learning Population-level Representations of Intracranial Activity

### v1

@@ 7 Discussion
We presented a self-supervised learning scheme for learning effective representations of intracranial activity from temporal embeddings. We find that pretraining the PopT results in better channel efficiency at fine-tuning time. This can reduce the number of electrodes needed in future experiments, which is critical for an invasive recording modality such as sEEG. We showed that self-supervised pretraining imbues our model with knowledge of spatial relationships between these embeddings and improved downstream decoding that scales with the number of available channels. As an aside, we note that the tasks we evaluate necessitate wide coverage of the brain. This is evidenced by the fact that performance scales with the number of input channels. With future collection of high quality intracranial data, we can continue scaling PopT and uncover exciting new data-driven findings for neuroscience.
By decoupling temporal and spatial feature extraction, we are able to leverage existing temporal embeddings to learn spatiotemporal representations efficiently and with a smaller number of parameters. Our approach also leaves open the possibility for independent improvement in temporal modeling. If future approaches introduce better time-series representations, are approach will be able to incorporate these advantages directly. Finally, we note that our method can serve more generally as a representation learning approach for any ensemble of sparsely distributed time-series data channels.
Limitations and Future Work As far as we know, no large public sEEG dataset that are of the same level of quality as ours (2048 Hz sampling rate, aligned electrode coordiantes, multimodal stimulus) are available, so direct comparison with existing approaches is difficult. Additionally, existing sEEG test datasets that have been used by existing deep learning models [21] focus on the artifact and seizure detection tasks [41], which are less interesting at a network-level due to the dependence on human labeling while looking at the time-series sEEG data [42].
Given the high sampling rate of our sEEG data (10x of prior work [21, 25]), training an end-to-end spatio-temporal model on our data would not have been computationally feasible, lending to the benefits of learning spatial representations on top of learned temporal embeddings. With the development and acquisition of compute resources, it would be a valuable future work to compare our approach with end-to-end approaches.
@@ 8 Conclusion
We introduced a pretraining method for learning representations of arbitrary ensembles of intracranial electrodes. We showed that our pretraining produced considerable improvements in downstream decoding, that would not have been possible without the knowledge of spatial relationships learned during the self-supervised pretraining stage. We showed that this scheme produces interpretable weights from which connectivity maps and candidate functional brain regions can be read. Finally, we release the pretrained weights for our PopT with BrainBERT inputs as well as our code for plug-and-play pretraining with any temporal embedding (see attached supplemental materials).
@@ 9 Acknowledgements
We would like to thank our funding sources: Caltech Chen Institute, Caltech Carver Mead New Adventures Fund. This work was also supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the NSF awards 2124052 and 2123818, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Artificial Social Intelligence for Successful Teams (ASIST) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0014, NIH grant R01EY026025, and the Office of Naval Research under award number N00014-20-1-2589 and award number N00014-20-1-2643. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We would also like to thank Xuefei (Julie) Wang for helpful comments and discussion.
@@ Appendix A Architectures and training
Pretrained PopT The core Population Transformer consists of a transformer encoder stack with 6 layers, 8 heads. All layers ($N=6$) in the encoder stack are set with the following parameters: $d_{h}=512$, $H=8$, and $p_{\text{dropout}}=0.1$. We pretrain the PopT model with the LAMB optimizer [43] ($lr=1e-4$), with a batch size of $n_{\text{batch}}=256$, and train/val/test split of 0.98, 0.01, 0.01 of the data. We pretrain for 500,000 steps, and record the validation performance every 1,000 steps. Downstream evaluation takes place on the weights with the best validation performance. We use the intermediate representation at the [CLS] token $d_{h}=512$ and put a linear layer that outputs to $d_{out}=1$ for fine-tuning on downstream tasks. These parameters for pretraining were the same for any PopT that needed to be pretrained (hold-one-out subject, subject subsets, ablation studies).
@@
Linear The linear baseline consists of a single linear layer that outputs to $d_{out}=1$. The inputs are flattened and concatenated BrainBERT embeddings $d_{emb}=756$ from a subset of channels $S\subset C$. Thus, the full input dimension is $d_{input}=d_{emb}*|S|$.
@@
Downstream Training For both PopT models, we train with these parameters: AdamW optimizer [44], $lr=5e^{-4}$ where transformer weights are scaled down by a factor of 10 ($lr_{t}=5e^{-5}$), $n_{batch}=256$, a Ramp Up scheduler [45] with warmup 0.025 and Step LR gamma 0.99, reducing 100 times within the 2000 total steps that we train for. For Linear and DeepNN models, we train with these parameters: AdamW optimizer [44], $lr=5e^{-4}$, $n_{batch}=256$, a Ramp Up scheduler [45] with warmup 0.025 and Step LR gamma 0.95, reducing 25 times within the 17,000 total steps we train for. For all downstream decoding, we use a fixed train/val/test split of 0.8, 0.1, 0.1 of the data.
Compute Resources To run all our experiments (data processing, pretraining, evaluations, interpretability), one only needs 1 NVIDIA Titan RTXs (24GB GPU Ram) and up to 80 CPU cores (700GB memory) if running sequentially. Pretraining PopT takes 4 days on 1 GPU. Our downstream evaluations take a few minutes to run each. For the purposes of data processing and gathering all the results in the paper, we parallelized the experiments on roughly 8 GPUs and 80 CPU cores.
Appendix B Decoding tasks
We follow the same task specification as in Wang et al. [8], with the modification that the pitch and volume examples are determined by percentile (see below) rather than standard deviation in order to obtain balanced classes.
@@ Appendix B Decoding tasks
Speech vs non-speech Positive examples are intervals of brain activity that correspond with dialogue being spoken in the stimuli movie. Negative examples are intervals of activity from 1s periods during which no speech is occurring in the movie.
Sentence onset Negative examples are as before. Positive examples are intervals of brain activity that correspond with hearing the first word of a sentence.
Appendix C Data
@@
Table 3: Subject statistics Subjects used in PopT training, and held-out downstream evaluation. Table taken from [8]. The number of uncorrupted, electrodes that can be Laplacian re-referenced are shown in the second column The average amount of recording data per subject is 4.3 (hrs).
Appendix D Interpretation Methods
Connectivity analysis We start with a pretrained PopT. To test a particular channel’s contribution to connectivity, we mask it with all zeros. Then, we consider the remaining unmasked channels and ask, how much increase do we see in the pretraining channel-wise loss? Recall that this objective is to determine whether or not a channel has had its inputs swapped with random activity. If the increase in loss is large, then we infer that the masked channel provided important context for this task. Using this delta as a measure for connectivity, we can then average across regions, as provided by the Desikan-Killiany atlas [46] and produce a plot using mne-connectivity [47].
Scaled Attention Weight First, we obtain an attention weight matrix across all trials which includes weights between all tokens. Then, we perform attention rollout [38] across layers to obtain the contributions of each input channel by the last layer. We take the resulting last layer of rollout weights for all channels, where the target is the [CLS] token, normalize within subject, and scale by ROC AUC to obtain the Scaled Attention Weight per channel. Finally, we plot the 0.75 percentile weight per region, as mapped by the Destrieux atlas [39] using Nilearn [48].
Appendix E Connectivity
Figure 9: Full connectivity for our 7 test subjects. We compare between traditional connectivity analaysis performed via coherence (top row in each section) and the analysis based on our PopT pretrained weights (bottom row in each section). We note that our analysis usually recovers the strongest points of connectivity fromt the traditional analysis. Coherence was computed using scikit-learn’s [49] signal.coherence.
Appendix F Frozen scaling
@@
Appendix G Frozen ablation

### v2

@@ 7 Discussion
We presented a self-supervised scheme for learning effective joint representations of neural activity from temporal embeddings. Our approach improves decoding and reduces the samples required to learn downstream tasks, which is especially critical for neural data modalities given patient constraints. A key aspect of our approach is the fact that we focus on spatial aggregation of existing channel embeddings, rather than training a large end-to-end model. By decoupling temporal and spatial feature extraction, we are able to leverage existing temporal embeddings to learn spatiotemporal representations efficiently and with a smaller number of parameters. This makes our model available for use in low compute-resource settings. Furthermore, this separation of considerations opens up the possibility for future independent improvement in temporal modeling, whether that be from a domain specific model or a more general time-series encoder. The generality of this approach allowed us to train on two very different neural modalities: scalp EEG and invasive iEEG. Our success in these domains suggest that this approach could even be extended to settings outside of neuroscience that also contend with sparsely and variably distributed time-series data channels, as is often the case with geophysical or climate data.
Limitations and Future Work We propose an effective strategy for aggregating signals with meaningful spatial coordinates, but it remains to be seen how we can generalize this further to aggregations without such coordinates. Individual brains are also highly variable, so improving this positional encoding scheme could provide an even higher resolution latent space of each individual’s brain. Future work could experiment with automatic functional identification for each channel, such as that explored in neural spiking data (Azabou et al., 2024), but it is currently unclear how to do so with neural recordings that have lower SNR.
@@ 8 Conclusion
We introduced a pretraining method for learning representations of arbitrary ensembles of intracranial electrodes. We showed that our pretraining produced considerable improvements in downstream decoding and efficiency, that would not have been possible without the knowledge of spatial relationships learned during the self-supervised pretraining stage. These benefits were found across data modalities, decoding tasks, and temporal encoders used, speaking to the generality of our approach. We further showed that this scheme produces interpretable weights from which connectivity maps and candidate functional brain regions can be read. Finally, we release the pretrained weights for our PopT with BrainBERT inputs as well as our code for plug-and-play pretraining with any temporal embedding.
@@ 9 Acknowledgements
We would like to thank our funding sources: Caltech Chen Institute, Caltech Carver Mead New Adventures Fund. This work was also supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the NSF awards 2124052 and 2123818, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Artificial Social Intelligence for Successful Teams (ASIST) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0014, NIH grant R01EY026025, and the Office of Naval Research under award number N00014-20-1-2589 and award number N00014-20-1-2643. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We would also like to thank Xuefei Wang and Amy Li for helpful comments and discussion.
@@ Appendix A Architectures and training
Pretrained PopT The core Population Transformer consists of a transformer encoder stack with 6 layers, 8 heads. All layers ($N=6$) in the encoder stack are set with the following parameters: $d_{h}=512$, $H=8$, and $p_{\text{dropout}}=0.1$. We pretrain the PopT model with the LAMB optimizer (You et al., 2019) ($lr=1e-4$), with a batch size of $n_{\text{batch}}=256$, and train/val/test split of 0.98, 0.01, 0.01 of the data. We pretrain for 500,000 steps, and record the validation performance every 1,000 steps. Downstream evaluation takes place on the weights with the best validation performance. We use the intermediate representation at the [CLS] token $d_{h}=512$ and put a linear layer that outputs to $d_{out}=1$ for fine-tuning on downstream tasks. These parameters for pretraining were the same for any PopT that needed to be pretrained (hold-one-out subject, subject subsets, ablation studies).
@@
Linear The linear baseline consists of a single linear layer that outputs to $d_{out}=1$. The inputs are flattened and concatenated BrainBERT embeddings $d_{emb}=756$, TOTEM embeddings $d_{emb}=64$, Chronos embeddings $d_{emb}=512$, or TS2Vec embeddings $d_{emb}=320$ from a subset of channels $S\subset C$. Thus, the full input dimension is $d_{input}=d_{emb}*|S|$.
@@
Downstream Training For both PopT models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=5e^{-4}$ where transformer weights are scaled down by a factor of 10 ($lr_{t}=5e^{-5}$), $n_{batch}=256$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.99, reducing 100 times within the 2000 total steps that we train for. For Linear and DeepNN models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=5e^{-4}$, $n_{batch}=256$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.95, reducing 25 times within the 17,000 total steps we train for. For all downstream decoding, we use a fixed train/val/test split of 0.8, 0.1, 0.1 of the data.
Compute Resources To run all our experiments (data processing, pretraining, evaluations, interpretability), one only needs 1 NVIDIA Titan RTXs (24GB GPU Ram). Pretraining PopT takes 2 days on 1 GPU. Our downstream evaluations take a few minutes to run each. For the purposes of data processing and gathering all the results in the paper, we parallelized the experiments on 8 GPUs.
Appendix B Model and Compute Requirements
e5 e50 e90
PopT 20M
Deep NN 3M 20M 36M
Linear 3.8k 38k 69k
Brant (Zhang et al., 2024) 500M
LaBraM (Jiang et al., 2024) 350M
Table 4: Parameter counts. Since PopT takes existing temporal embeddings as input, the number of parameters that must be trained is an order of magnitude less than recent end-to-end approaches.
GPU count GPU type Time to train TFLOPS
PopT 1 NVIDIA TITAN RTX (24GB) 2 days 2.1M
Brant (Zhang et al., 2024) 4 NVIDIA Tesla A100 (80G) 2.8 days 18.8M
LaBraM (Jiang et al., 2024) 8 NVIDIA Tesla A800 (40G)
Table 5: Pretraining compute requirements Based on published train times (none were given for LaBraM) it is evident that PopT has smaller hardware and shorter training time requirements.
Appendix C Decoding tasks
We follow the same task specification as in Wang et al. (2022), with the modification that the pitch and volume examples are determined by percentile (see below) rather than standard deviation in order to obtain balanced classes.
@@ Appendix C Decoding tasks
Sentence onset Negative examples are intervals of activity from 1s periods during which no speech is occurring in the movie. Positive examples are intervals of brain activity that correspond with hearing the first word of a sentence.
Speech vs. Non-speech Negative examples are as before. Positive examples are intervals of brain activity that correspond with dialogue being spoken in the stimuli movie.
Appendix D Data
@@
Table 6: Subject statistics Subjects used in PopT training, and held-out downstream evaluation. Table taken from Wang et al. (2022). The second column shows the number of uncorrupted, electrodes that can be Laplacian re-referenced. The average amount of recording data per subject is 4.3 (hrs).
Appendix E Interpretation Methods
Connectivity analysis We start with a pretrained PopT. To test a particular channel’s contribution to connectivity, we mask it with all zeros. Then, we consider the remaining unmasked channels and ask, how much increase do we see in the pretraining channel-wise loss? Recall that this objective is to determine whether or not a channel has had its inputs swapped with random activity. If the increase in loss is large, then we infer that the masked channel provided important context for this task. Using this delta as a measure for connectivity, we then average across the Desikan-Killiany atlas regions (Alexander et al., 2019) and produce a plot using mne-connectivity (Gramfort et al., 2013).
Scaled Attention Weight First, we obtain an attention weight matrix across all trials which includes weights between all tokens. Then, we perform attention rollout (Abnar & Zuidema, 2020) across layers to obtain the contributions of each input channel by the last layer. We take the resulting last layer of rollout weights for all channels, where the target is the [CLS] token, normalize within subject, and scale by ROC AUC to obtain the Scaled Attention Weight per channel. Finally, we plot the 0.75 percentile weight per region, as mapped by the Destrieux atlas (Destrieux et al., 2010) using Nilearn (contributors, ).
Appendix F Connectivity
Figure 10: Full connectivity for our 7 test subjects. We compare between traditional connectivity analaysis performed via coherence (top row in each section) and the analysis based on our PopT pretrained weights (bottom row in each section). We note that our analysis usually recovers the strongest points of connectivity fromt the traditional analysis. Coherence was computed using scikit-learn’s (Pedregosa et al., 2011) signal.coherence.
Appendix G Frozen scaling
@@
Appendix H Frozen ablation